Two cards are drawn from a deck. How many possible values can each of the following variables take?

  • Sum of the numbers on the cards:
  • Number of times both cards are black:
  • Number of times both cards are 7s:
  • Number of times the first card is six and the second card is red:
  • Number of times the first card is face card and the second card is not a face card:
1
Expert's answer
2022-02-10T17:50:10-0500

"a)"

In a deck of 52 cards, there are cards numbered from 1 to 13.

When two cards are drawn, least sum obtained is 2 (=1+1) and greatest sum obtained is 26 (=13+13).


"b)"

There are 26 black cards.

Since two cards are drawn, then we have "\binom{26}{2}={26\times25\over2}=325"

Therefore, there are 325 ways in which both cards are black


"c)"

There are four 7's and the number of ways in which both cards are 7's is "\binom{4}{2}=2\times3=6" ways


"d)"

There are four 6's and 26 red cards. The number of  the first card is six and the second card is red is,

"\binom{4}{1}\times\binom{26}{1}=4\times26=104" times

"e)"

There are 12 face cards and 40 non-faced cards.

So, number of times the first card is face card and the second card is not a face card from deck of 52 cards is "\binom{12}{1}\times \binom{40}{1} =12\times40=480" times.